\begin{thebibliography}{69}

\bibitem[Ahmad \emph{et~al.}(2024)Ahmad, Sadhasivam, Lisa, Lombardo, Emil,
  Zaki, {Van Westen}, Fadel and Tanyas]{Ahmad2023}
Ahmad, S.~M., Sadhasivam, N., Lisa, M., Lombardo, L., Emil, M.~K., Zaki, A.,
  {Van Westen}, C.~J., Fadel, I. and Tanyas, H. (2024) {Standing on the
  shoulder of a giant landslide: A six-year long InSAR look at a slow-moving
  hillslope in the western Karakoram}.
\newblock \emph{Geomorphology} \textbf{444}, 108959.

\bibitem[Behrens \emph{et~al.}(2004)Behrens, Lopes and
  Gamerman]{Behrens.etal:2004}
Behrens, C., Lopes, H. and Gamerman, D. (2004) {Bayesian analysis of extreme
  events with threshold estimation}.
\newblock \emph{{Statistical Modelling}} \textbf{4}, 227--244.

\bibitem[Beirlant \emph{et~al.}(2019)Beirlant, Kijko, Reynkens and
  Einmahl]{Beirlant2019Estimating}
Beirlant, J., Kijko, A., Reynkens, T. and Einmahl, J. H.~J. (2019) Estimating
  the maximum possible earthquake magnitude using extreme value methodology:
  the groningen case.
\newblock \emph{Natural Hazards} \textbf{98}, 1091--1113.

\bibitem[Besag(1974)]{besag1974spatial}
Besag, J. (1974) Spatial interaction and the statistical analysis of lattice
  systems.
\newblock \emph{Journal of the {R}oyal {S}tatistical {S}ociety: {S}eries B
  ({M}ethodological)} \textbf{36}, 192--225.

\bibitem[Van~den Bout \emph{et~al.}(2023)Van~den Bout, Jetten, van Westen and
  Lombardo]{van2023breakthrough}
Van~den Bout, B., Jetten, V., van Westen, C.~J. and Lombardo, L. (2023) A
  breakthrough in fast flood simulation.
\newblock \emph{Environmental Modelling \& Software} \textbf{168}, 105787.

\bibitem[Carreau and Bengio(2009)]{Carreau.Bengio:2009}
Carreau, J. and Bengio, Y. (2009) {A hybrid Pareto model for asymmetric
  fat-tailed data: the univariate case}.
\newblock \emph{{Extremes}} \textbf{12}, 53--76.

\bibitem[Castro-Camilo \emph{et~al.}(2019)Castro-Camilo, Huser and
  Rue]{castro2019spliced}
Castro-Camilo, D., Huser, R. and Rue, H. (2019) A spliced {G}amma-generalized
  {P}areto model for short-term extreme wind speed probabilistic forecasting.
\newblock \emph{Journal of Agricultural, Biological and Environmental
  Statistics} \textbf{24}, 517--534.

\bibitem[Cisneros \emph{et~al.}(2024)Cisneros, Richards, Dahal, Lombardo,  and
  Huser]{cisneros2024deep}
Cisneros, D., Richards, J., Dahal, A., Lombardo, L.,  and Huser, R. (2024) Deep
  graphical regression models for jointly moderate and extreme australian
  wildfires.
\newblock \emph{Spatial Statistics} \textbf{59}, 100811.

\bibitem[Cooley \emph{et~al.}(2007)Cooley, Nychka and
  Naveau]{cooley2007bayesian}
Cooley, D., Nychka, D. and Naveau, P. (2007) Bayesian spatial modeling of
  extreme precipitation return levels.
\newblock \emph{Journal of the American Statistical Association} \textbf{102},
  824--840.

\bibitem[Dahal \emph{et~al.}(2023)Dahal, Castro-Cruz, Tanya{\c{s}}, Fadel, Mai,
  van~der Meijde, van Westen, Huser and Lombardo]{dahal2023ground}
Dahal, A., Castro-Cruz, D.~A., Tanya{\c{s}}, H., Fadel, I., Mai, P.~M., van~der
  Meijde, M., van Westen, C., Huser, R. and Lombardo, L. (2023) From ground
  motion simulations to landslide occurrence prediction.
\newblock \emph{Geomorphology} \textbf{441}, 108898.

\bibitem[Dahal \emph{et~al.}(2024)Dahal, Huser and Lombardo]{dahal2024junction}
Dahal, A., Huser, R. and Lombardo, L. (2024) At the junction between deep
  learning and statistics of extremes: formalizing the landslide hazard
  definition.
\newblock arXiv preprint 2401.14210.

\bibitem[Darzi \emph{et~al.}(2023)Darzi, Hrafnkelsson and
  Halldorsson]{darzi2023Bayesian}
Darzi, A., Hrafnkelsson, B. and Halldorsson, B. (2023) {Bayesian modelling in
  engineering seismology: spatial earthquake magnitude model}.
\newblock In \emph{{Bayesian Modelling in Engineering Seismology: Spatial
  Earthquake Magnitude Model}}, ed. B.~Hrafnkelsson, pp. 171--192. {Springer}.

\bibitem[Davison and Gholamrezaee(2012)]{Davison.Gholamrezaee:2012}
Davison, A.~C. and Gholamrezaee, M.~M. (2012) {Geostatistics of extremes}.
\newblock \emph{{Proceedings of the Royal Society A: Mathematical, Physical \&
  Engineering Sciences}} \textbf{468}, 581--608.

\bibitem[Davison and Huser(2015)]{Davison.Huser:2015}
Davison, A.~C. and Huser, R. (2015) {Statistics of extremes}.
\newblock \emph{Annual Review of Statistics and its Application} \textbf{2},
  203--235.

\bibitem[Davison and Smith(1990)]{Davison.Smith:1990}
Davison, A.~C. and Smith, R.~L. (1990) {Models for exceedances over high
  thresholds (with discussion)}.
\newblock \emph{{Journal of the Royal Statistical Society: Series B
  (Statistical Methodology)}} \textbf{52}, 393--442.

\bibitem[Dawkins and Stephenson(2018)]{Dawkins2018}
Dawkins, L.~C. and Stephenson, D.~B. (2018) Quantification of extremal
  dependence in spatial natural hazard footprints: independence of windstorm
  gust speeds and its impact on aggregate losses.
\newblock \emph{Natural Hazards and Earth System Sciences} \textbf{18},
  2933--2949.

\bibitem[Di~Napoli \emph{et~al.}(2023)Di~Napoli, Tanyas, Castro-Camilo,
  Calcaterra, Cevasco, Di~Martire, Pepe, Brandolini and Lombardo]{DiNapoli2023}
Di~Napoli, M., Tanyas, H., Castro-Camilo, D., Calcaterra, D., Cevasco, A.,
  Di~Martire, D., Pepe, G., Brandolini, P. and Lombardo, L. (2023) On the
  estimation of landslide intensity, hazard and density via data-driven models.
\newblock \emph{Natural Hazards} .

\bibitem[Dutfoy and Senfaute(2022)]{dutfoy2022randomized}
Dutfoy, A. and Senfaute, G. (2022) The randomized gutenberg--richter model: a
  recurrence model based on extreme value theory---impacts on probabilistic
  seismic hazard analyses and comparison with the standard approach.
\newblock \emph{Bulletin of Earthquake Engineering} \textbf{20}, 6349--6376.

\bibitem[de~Fondeville and Davison(2018)]{deFondevilleDavison18}
de~Fondeville, R. and Davison, A.~C. (2018) High-dimensional
  peaks-over-threshold inference.
\newblock \emph{Biometrika} \textbf{105}, 575--592.

\bibitem[Formetta \emph{et~al.}(2021)Formetta, Over and
  Stewart]{formetta2021assessment}
Formetta, G., Over, T. and Stewart, E. (2021) {Assessment of peak flow scaling
  and its effect on flood quantile estimation in the United Kingdom}.
\newblock \emph{Water Resources Research} \textbf{57}, e2020WR028076.

\bibitem[Guzzetti \emph{et~al.}(1999)Guzzetti, Carrara, Cardinali and
  Reichenbach]{guzzetti1999landslidehazard}
Guzzetti, F., Carrara, A., Cardinali, M. and Reichenbach, P. (1999) Landslide
  hazard evaluation: A review of current techniques and their application in a
  multi-scale study, central {Italy}.
\newblock \emph{Geomorphology} \textbf{31}, 181--216.

\bibitem[Guzzetti \emph{et~al.}(2007)Guzzetti, Peruccacci, M. and
  Stark]{guzzetti2007rainfall}
Guzzetti, F., Peruccacci, S., M., R. and Stark, C.~P. (2007) Rainfall
  thresholds for the initiation of landslides in central and southern europe.
\newblock \emph{Meteorology and Atmospheric Physics} \textbf{98}, 239--267.

\bibitem[He \emph{et~al.}(2023)He, Tanyas, Chang, Hu, Luo and
  Lombardo]{he2023modelling}
He, K., Tanyas, H., Chang, L., Hu, X., Luo, G. and Lombardo, L. (2023)
  {Modelling InSAR-derived hillslope velocities with multivariate statistics: A
  first attempt to generate interpretable predictions}.
\newblock \emph{Remote Sensing of Environment} \textbf{289}, 113518.

\bibitem[Hough(2014)]{hough2014earthquake}
Hough, S.~E. (2014) {Earthquake intensity distributions: A new view}.
\newblock \emph{Bulletin of earthquake engineering} \textbf{12}, 135--155.

\bibitem[Hungr(2018)]{hungr2018some}
Hungr, O. (2018) Some methods of landslide hazard intensity mapping.
\newblock In \emph{Landslide risk assessment}, pp. 215--226. Routledge.

\bibitem[Huser and Davison(2014)]{huser2014space}
Huser, R. and Davison, A.~C. (2014) Space-time modelling of extreme events.
\newblock \emph{Journal of the Royal Statistical Society: Series B}
  \textbf{76}, 439--461.

\bibitem[Huser \emph{et~al.}(2021)Huser, Opitz and Thibaud]{huser2021maxid}
Huser, R., Opitz, T. and Thibaud, E. (2021) Max-infinitely divisible models and
  inference for spatial extremes.
\newblock \emph{Scandinavian Journal of Statistics} \textbf{48}, 321--348.

\bibitem[Huser \emph{et~al.}(2024)Huser, Opitz and
  Wadsworth]{huser2024modeling}
Huser, R., Opitz, T. and Wadsworth, J. (2024) Modeling of spatial extremes in
  environmental data science: Time to move away from max-stable processes.
\newblock arXiv preprint 2401.17430.

\bibitem[Illian \emph{et~al.}(2008)Illian, Penttinen, Helga and
  Stoyan]{illian2008}
Illian, J., Penttinen, A., Helga, S. and Stoyan, D. (2008) \emph{Statistical
  Analysis and Modelling of Spatial Point Patterns}.
\newblock John Wiley \& Sons, Ltd.

\bibitem[Jaboyedoff \emph{et~al.}(2020)Jaboyedoff, Carrea, Derron, Oppikofer,
  Penna and Rudaz]{jaboyedoff2020review}
Jaboyedoff, M., Carrea, D., Derron, M.-H., Oppikofer, T., Penna, I.~M. and
  Rudaz, B. (2020) A review of methods used to estimate initial landslide
  failure surface depths and volumes.
\newblock \emph{Engineering Geology} \textbf{267}, 105478.

\bibitem[Karlsson \emph{et~al.}(2017)Karlsson, Kalantari, M{\"o}rtberg,
  Olofsson and Lyon]{karlsson2017natural}
Karlsson, C.~S., Kalantari, Z., M{\"o}rtberg, U., Olofsson, B. and Lyon, S.~W.
  (2017) Natural hazard susceptibility assessment for road planning using
  spatial multi-criteria analysis.
\newblock \emph{Environmental management} \textbf{60}, 823--851.

\bibitem[Katz \emph{et~al.}(2002)Katz, Parlange and Naveau]{katz2002statistics}
Katz, R.~W., Parlange, M.~B. and Naveau, P. (2002) Statistics of extremes in
  hydrology.
\newblock \emph{Advances in Water Resources} \textbf{25}, 1287--1304.

\bibitem[Kiriliouk \emph{et~al.}(2019)Kiriliouk, Rootz\'en, Segers and
  Wadsworth]{Kiriliouk2019peaks}
Kiriliouk, A., Rootz\'en, H., Segers, J. and Wadsworth, J.~L. (2019) Peaks over
  thresholds modeling with multivariate generalized {P}areto distributions.
\newblock \emph{Technometrics} \textbf{61}, 123--135.

\bibitem[Koh \emph{et~al.}(2023)Koh, Pimont, Dupuy and
  Opitz]{koh2021spatiotemporal}
Koh, J., Pimont, F., Dupuy, J.-L. and Opitz, T. (2023) Spatiotemporal wildfire
  modeling through point processes with moderate and extreme marks.
\newblock \emph{{Annals of Applied Statistics,}} \textbf{17}, 560--582.

\bibitem[Lari \emph{et~al.}(2014)Lari, Frattini and
  Crosta]{lari2014probabilistic}
Lari, S., Frattini, P. and Crosta, G. (2014) A probabilistic approach for
  landslide hazard analysis.
\newblock \emph{Engineering geology} \textbf{182}, 3--14.

\bibitem[Lombardo \emph{et~al.}(2019)Lombardo, Bakka, Tanyas, van Westen, Mai
  and Huser]{lombardo2019geostatistical}
Lombardo, L., Bakka, H., Tanyas, H., van Westen, C., Mai, P.~M. and Huser, R.
  (2019) Geostatistical modeling to capture seismic-shaking patterns from
  earthquake-induced landslides.
\newblock \emph{Journal of {G}eophysical {R}esearch: {E}arth {S}urface}
  \textbf{124}, 1958--1980.

\bibitem[Lombardo \emph{et~al.}(2020)Lombardo, Opitz, Ardizzone, Guzzetti and
  Huser]{lombardo2020space}
Lombardo, L., Opitz, T., Ardizzone, F., Guzzetti, F. and Huser, R. (2020)
  Space-time landslide predictive modelling.
\newblock \emph{{Earth-Science Reviews}} \textbf{209}, 103318.

\bibitem[Lombardo \emph{et~al.}(2018)Lombardo, Opitz and
  Huser]{lombardo2018point}
Lombardo, L., Opitz, T. and Huser, R. (2018) Point process-based modeling of
  multiple debris flow landslides using {INLA}: an application to the 2009
  {M}essina disaster.
\newblock \emph{{Stochastic Environmental Research and Risk Assessment}}
  \textbf{32}, 2179--2198.

\bibitem[Lombardo \emph{et~al.}(2021)Lombardo, Tanyas, Huser, Guzzetti and
  Castro-Camilo]{lombardo2021landslide}
Lombardo, L., Tanyas, H., Huser, R., Guzzetti, F. and Castro-Camilo, D. (2021)
  Landslide size matters: A new data-driven, spatial prototype.
\newblock \emph{Engineering {G}eology} \textbf{293}, 106288.

\bibitem[Mazzanti \emph{et~al.}(2015)Mazzanti, Bozzano, Cipriani and
  Prestininzi]{mazzanti2015new}
Mazzanti, P., Bozzano, F., Cipriani, I. and Prestininzi, A. (2015) {New
  insights into the temporal prediction of landslides by a terrestrial SAR
  interferometry monitoring case study}.
\newblock \emph{Landslides} \textbf{12}, 55--68.

\bibitem[Mitchell(1993)]{mitchell1993natural}
Mitchell, J.~K. (1993) Natural hazard predictions and responses in very large
  cities.
\newblock \emph{Prediction and Perception of Natural Hazards. Dordrecht, The
  Netherlands: Kluwer Academic Publishers} pp. 29--37.

\bibitem[M{\o}ller \emph{et~al.}(1998)M{\o}ller, Syversveen and
  Waagepetersen]{moller1998log}
M{\o}ller, J., Syversveen, A.~R. and Waagepetersen, R.~P. (1998) Log {G}aussian
  {C}ox processes.
\newblock \emph{Scandinavian {J}ournal of {S}tatistics} \textbf{25}, 451--482.

\bibitem[Murphy and O'brien(1977)]{murphy1977correlation}
Murphy, J.~u. and O'brien, L. (1977) The correlation of peak ground
  acceleration amplitude with seismic intensity and other physical parameters.
\newblock \emph{Bulletin of the Seismological Society of America} \textbf{67},
  877--915.

\bibitem[Naveau \emph{et~al.}(2016)Naveau, Huser, Ribereau and
  Hannart]{Naveau.etal:2016}
Naveau, P., Huser, R., Ribereau, P. and Hannart, A. (2016) {Modeling jointly
  low, moderate, and heavy rainfall intensities without a threshold selection}.
\newblock \emph{Water Resources Research} \textbf{52}, 2753--2769.

\bibitem[Nicu \emph{et~al.}(2022)Nicu, Elia, Rubensdotter, Tanyas and
  Lombardo]{nicu2022multi}
Nicu, I.~C., Elia, L., Rubensdotter, L., Tanyas, H. and Lombardo, L. (2022)
  {Multi-hazard susceptibility mapping of cryospheric hazards in a high-Arctic
  environment: Svalbard Archipelago}.
\newblock \emph{Earth System Science Data Discussions} \textbf{2022}, 1--26.

\bibitem[Opitz(2016)]{opitz2016modeling}
Opitz, T. (2016) Modeling asymptotically independent spatial extremes based on
  {L}aplace random fields.
\newblock \emph{Spatial Statistics} \textbf{16}, 1--18.

\bibitem[Opitz \emph{et~al.}(2018)Opitz, Huser, Bakka and Rue]{opitz2018inla}
Opitz, T., Huser, R., Bakka, H. and Rue, H. (2018) {INLA} goes extreme:
  Bayesian tail regression for the estimation of high spatio-temporal
  quantiles.
\newblock \emph{Extremes} \textbf{21}, 441--462.

\bibitem[Papastathopoulos and Tawn(2013)]{Papastathopoulos.Tawn:2013}
Papastathopoulos, I. and Tawn, J.~A. (2013) {Extended generalised Pareto models
  for tail estimation}.
\newblock \emph{{Journal of Statistical Planning and Inference}} \textbf{143},
  131--143.

\bibitem[Peng \emph{et~al.}(2005)Peng, Schoenberg and Woods]{peng2005space}
Peng, R.~D., Schoenberg, F.~P. and Woods, J.~A. (2005) A space--time
  conditional intensity model for evaluating a wildfire hazard index.
\newblock \emph{Journal of the American Statistical Association} \textbf{100},
  26--35.

\bibitem[Pimont \emph{et~al.}(2021)Pimont, Fargeon, Opitz, Ruffault, Barbero,
  Martin-StPaul, Rigolot, Rivi{\`e}re and Dupuy]{pimont2021prediction}
Pimont, F., Fargeon, H., Opitz, T., Ruffault, J., Barbero, R., Martin-StPaul,
  N., Rigolot, E., Rivi{\`e}re, M. and Dupuy, J.-L. (2021) Prediction of
  regional wildfire activity in the probabilistic {B}ayesian framework of
  firelihood.
\newblock \emph{Ecological {A}pplications} \textbf{31}, e02316.

\bibitem[Pudasaini and Krautblatter(2021)]{pudasaini2021mechanics}
Pudasaini, S.~P. and Krautblatter, M. (2021) The mechanics of landslide
  mobility with erosion.
\newblock \emph{Nature communications} \textbf{12}, 6793.

\bibitem[Pudasaini and Krautblatter(2022)]{pudasaini2022landslide}
Pudasaini, S.~P. and Krautblatter, M. (2022) The landslide velocity.
\newblock \emph{Earth Surface Dynamics} \textbf{10}, 165--189.

\bibitem[Rana \emph{et~al.}(2023)Rana, Bhuyan, Ferrer, Cotton, Ozturk, Catani
  and Malik]{rana2023landslide}
Rana, K., Bhuyan, K., Ferrer, J.~V., Cotton, F., Ozturk, U., Catani, F. and
  Malik, N. (2023) {Landslide Topology Uncovers Failure Movements}.
\newblock \emph{arXiv preprint arXiv:2310.09631} .

\bibitem[Risser and Wehner(2017)]{Risser2017Attributable}
Risser, M.~D. and Wehner, M.~F. (2017) Attributable human-induced changes in
  the likelihood and magnitude of the observed extreme precipitation during
  {H}urricane {H}arvey.
\newblock \emph{Geophysical Research Letters} \textbf{28}, 12457--12464.

\bibitem[Scarrott and {MacDonald}(2012)]{Scarrott.MacDonald:2012}
Scarrott, C. and {MacDonald}, A. (2012) {A Review of Extreme Value Threshold
  Estimation And Uncertainty Quantification}.
\newblock \emph{{REVSTAT}} \textbf{10}, 33--60.

\bibitem[Stein(2021{a})]{Stein:2021a}
Stein, M.~L. (2021{a}) A parametric model for distributions with flexible
  behavior in both tails.
\newblock \emph{Environmetrics} \textbf{32}, e2658.

\bibitem[Stein(2021{b})]{Stein:2021b}
Stein, M.~L. (2021{b}) Parametric models for distributions when interest is in
  extremes with an application to daily temperature.
\newblock \emph{Extremes} \textbf{24}, 293---323.

\bibitem[Tang \emph{et~al.}(2014)Tang, Hu, Xu, Li, Yong and
  Wang]{tang2014novel}
Tang, H., Hu, X., Xu, C., Li, C., Yong, R. and Wang, L. (2014) A novel approach
  for determining landslide pushing force based on landslide-pile interactions.
\newblock \emph{Engineering Geology} \textbf{182}, 15--24.

\bibitem[Trugman \emph{et~al.}(2019)Trugman, Page, Minson and
  Cochran]{trugman2019peak}
Trugman, D.~T., Page, M.~T., Minson, S.~E. and Cochran, E.~S. (2019) {Peak
  ground displacement saturates exactly when expected: Implications for
  earthquake early warning}.
\newblock \emph{Journal of Geophysical Research: Solid Earth} \textbf{124},
  4642--4653.

\bibitem[Tseng \emph{et~al.}(2013)Tseng, Lin, Stark, Liu, Fei and
  Hsieh]{tseng2013application}
Tseng, C.-M., Lin, C.-W., Stark, C.~P., Liu, J.-K., Fei, L.-Y. and Hsieh, Y.-C.
  (2013) {Application of a multi-temporal, LiDAR-derived, digital terrain model
  in a landslide-volume estimation}.
\newblock \emph{Earth Surface Processes and Landforms} \textbf{38}, 1587--1601.

\bibitem[Varnes(1984)]{varnes1984}
Varnes, D.~J. (1984) \emph{Landslide Hazard Zonation: A Review of Principles
  and Practice}.
\newblock Natural Hazards. UNESCO, Paris.

\bibitem[Vikki \emph{et~al.}(2023)Vikki, Mitchell, Hegerl, Collins, Leach and
  Slingo]{thompson2023most}
Vikki, T., Mitchell, D., Hegerl, G.~C., Collins, M., Leach, N.~J. and Slingo,
  J.~M. (2023) The most at-risk regions in the world for high-impact heatwaves.
\newblock \emph{Nature Communications} \textbf{14}, 2152.

\bibitem[Vojtek and Vojtekov{\'a}(2016)]{vojtek2016flood}
Vojtek, M. and Vojtekov{\'a}, J. (2016) Flood hazard and flood risk assessment
  at the local spatial scale: a case study.
\newblock \emph{Geomatics, Natural Hazards and Risk} \textbf{7}, 1973--1992.

\bibitem[Winter and Tawn(2016)]{winter2016heatwaves}
Winter, H.~C. and Tawn, J.~A. (2016) Modelling heatwaves in central france: a
  case-study in extremal dependence.
\newblock \emph{Journal of the Royal Statistical Society: Series C}
  \textbf{65}, 345--365.

\bibitem[Yadav \emph{et~al.}(2021)Yadav, Huser and Opitz]{yadav2021spatial}
Yadav, R., Huser, R. and Opitz, T. (2021) Spatial hierarchical modeling of
  threshold exceedances using rate mixtures.
\newblock \emph{Environmetrics} \textbf{32}, e2662.

\bibitem[Yadav \emph{et~al.}(2022)Yadav, Huser and Opitz]{yadav2021flexible}
Yadav, R., Huser, R. and Opitz, T. (2022) A flexible {B}ayesian hierarchical
  modeling framework for spatially dependent peaks-over-threshold data.
\newblock \emph{Spatial Statistics,} \textbf{51}, 100672.

\bibitem[Yadav \emph{et~al.}(2023)Yadav, Huser, Opitz and
  Lombardo]{yadav2023joint}
Yadav, R., Huser, R., Opitz, T. and Lombardo, L. (2023) Joint modeling of
  landslide counts and sizes using spatial marked point processes with
  sub-asymptotic mark distributions.
\newblock \emph{Journal of the Royal Statistical Society Series C: Applied
  Statistics, 10.1093/jrsssc/qlad077} .

\bibitem[Zhang \emph{et~al.}(2023)Zhang, Risser, Wehner and
  O'Brien]{Zhang2023explaining}
Zhang, L., Risser, M.~D., Wehner, M.~F. and O'Brien, T.~A. (2023) Explaining
  the unexplainable: leveraging extremal dependence to characterize the 2021
  {P}acific {N}orthwest heatwave.
\newblock arXiv preprint 2307.03688.

\bibitem[Zhong \emph{et~al.}(2022)Zhong, Huser and Opitz]{Zhong.etal:2021}
Zhong, P., Huser, R. and Opitz, T. (2022) {Modeling non-stationary temperature
  maxima based on extremal dependence changing with event magnitude}.
\newblock \emph{{Annals of Applied Statistics}} \textbf{16}, 272--299.

\end{thebibliography}

\documentclass[11pt]{article}
\RequirePackage[OT1]{fontenc}
\usepackage{amsthm,amsmath,amssymb,amsfonts,bm,enumerate,xcolor,graphicx,natbib,color,url}
\RequirePackage[colorlinks,citecolor=blue,urlcolor=blue]{hyperref}
\usepackage{hyperref}
\usepackage{subcaption} 
\usepackage{ulem}
\usepackage{pifont}
\usepackage{textcmds}
\usepackage{booktabs}
\usepackage{multirow}
\usepackage{booktabs}
\usepackage{adjustbox}
\usepackage[usestackEOL]{stackengine}
\usepackage[toc,page]{appendix} 
\usepackage{algorithm,algorithmic}
\usepackage{lipsum}
\usepackage{setspace}

\usepackage{setspace}
\allowdisplaybreaks
\def\frac#1#2{{\textstyle{#1\over#2}}}

\DeclareSymbolFont{AMSb}{U}{msb}{m}{n}
\DeclareMathSymbol{\Natural}{\mathbin}{AMSb}{"4E}
\DeclareMathSymbol{\Integer}{\mathbin}{AMSb}{"5A}
\DeclareMathSymbol{\Real}{\mathbin}{AMSb}{"52}
\DeclareMathSymbol{\Rational}{\mathbin}{AMSb}{"51}
\DeclareMathSymbol{\Imaginary}{\mathbin}{AMSb}{"49}
\DeclareMathSymbol{\Complex}{\mathbin}{AMSb}{"43} 
\DeclareMathSymbol{\Disk}{\mathbin}{AMSb}{"44} 
\def\bi{\begin{itemize}}
\def\ei{\end{itemize}}
\def\bd{\begin{description}}
\def\ed{\end{description}}
\def\ben{\begin{enumerate}}
\def\een{\end{enumerate}}
\def\bv{\begin{verbatim}}
\def\ev{\end{verbatim}}

\def\2to{{\ {\buildrel 2\over \longrightarrow}\ }}

\def\I1ton{{$I_1,\ldots,I_n$}}
\def\X1ton{{$X_1,\ldots,X_n$}}
\def\Y1ton{{$Y_1,\ldots,Y_n$}}
\def\Z1ton{{$Z_1,\ldots,Z_n$}}
\def\R1ton{{$R_1,\ldots,R_n$}}
\def\e1ton{{$e_1,\ldots,e_n$}}
\def\t1ton{{$t_1,\ldots,t_n$}}
\def\x1ton{{$x_1,\ldots,x_n$}}
\def\y1ton{{$y_1,\ldots,y_n$}}
\def\z1ton{{$z_1,\ldots,z_n$}}
\newcommand{\blind}{1}

\begin{document}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\thispagestyle{empty}
\baselineskip=28pt
\vskip 5mm

\renewcommand{\thefootnote}{\fnsymbol{footnote}}

\begin{center} 
{\Large{\bf Statistics of extremes for natural hazards: landslides and earthquakes}}\footnotemark[2]
\end{center}

\baselineskip=12pt

\vskip 5mm

\renewcommand{\thefootnote}{\arabic{footnote}}

\if1\blind
{
\begin{center}
\large
Rishikesh Yadav$^1$, Luigi Lombardo$^2$, and Rapha\"el Huser$^3$
\end{center}
\renewcommand{\thefootnote}{\fnsymbol{footnote}} \footnotetext[2]{Extract of Chapter 27 (first 5 pages), to be published in the \emph{Handbook of Statistics of Extremes}, eds.\ Miguel de Carvalho, Rapha\"el Huser, Philippe Naveau and Brian Reich}
\renewcommand{\thefootnote}{\arabic{footnote}} \footnotetext[1]{
\baselineskip=10pt  Department of Decision Sciences, HEC Montr\'eal, Montr\'eal, Canada. E-mail: rishikesh.yadav@hec.ca}
\footnotetext[2]{
\baselineskip=10pt  ITC, Faculty of Geo-Information Science and Earth Observation, University of Twente, Enschede, The Netherlands. E-mail: l.lombardo@utwente.nl}
\footnotetext[3]{
\baselineskip=10pt Statistics Program, Computer, Electrical and Mathematical Sciences and Engineering (CEMSE) Division, King Abdullah University of Science and Technology (KAUST), Thuwal 23955-6900, Saudi Arabia. E-mail: raphael.huser@kaust.edu.sa}
\fi

\baselineskip=26pt
\vskip 2mm
\centerline{\today}
\vskip 4mm

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\baselineskip=14pt

In this chapter, we illustrate the use of split bulk--tail models and subasymptotic models motivated by extreme-value theory in the context of hazard assessment for earthquake-induced landslides. A spatial joint areal model is presented for modeling both landslides counts and landslide sizes, paying particular attention to extreme landslides, which are the most devastating ones.

\section{Introduction}
%\subsection{Background and motivation}
Statistics of extremes has been used extensively in climate science for the modeling of low-probability high-impact events, including natural hazards such as heavy precipitation \citep{katz2002statistics,cooley2007bayesian,huser2014space,deFondevilleDavison18}, extreme heatwaves \citep{Davison.Gholamrezaee:2012,winter2016heatwaves,Zhong.etal:2021,thompson2023most,Zhang2023explaining}, and strong windstorms or hurricanes \citep{opitz2016modeling,Risser2017Attributable,Dawkins2018,huser2021maxid}, among others; such extreme-value models are often developed with the ultimate ambition to improve the state-of-the-art in performing hazard and/or risk assessment and mitigation, and in attributing specific catastrophic extreme events to human influence under climate change. By contrast, the application of extreme-value theory (EVT) and statistics to the modeling and prediction of geophysical processes, such as devastating landslides or earthquakes, is sparser in the literature, perhaps due to the different type of data involved and the added modeling difficulties that come with it. Unlike climate data that are often measured at fixed locations (e.g., monitoring stations, or on a spatial grid with climate model outputs) and regular intervals (e.g., hourly or daily) over a period of time, the exact location and timing of landslides and earthquakes are typically unknown before they occur---they are thus often treated as random; moreover, the exact conditions and triggering mechanisms often differ for each event, which implies that the data are rarely replicated (unless some kind of spatio-temporal aggregation is used), making any extreme-value analysis more challenging, and that they typically require other kinds of specialized statistical models such as point processes \citep{moller1998log,illian2008}. Nevertheless, statistical methods based on EVT have still been used in data-driven geophysical sciences, e.g., for estimating the maximum earthquake magnitude \citep{Beirlant2019Estimating,darzi2023Bayesian}, for probabilistic seismic hazard analysis \citep{dutfoy2022randomized}, as well as for landslide hazard mapping over space \citep{yadav2023joint,dahal2024junction}; see also \cite{Kiriliouk2019peaks} who fit a multivariate extreme-value model to extreme rainfall data, in order to (indirectly) understand the probability of rainfall-induced landslides through the rainfall intensity--duration thresholds for landslide initiation established by \cite{guzzetti2007rainfall}. 

In this chapter, we focus on the spatial modeling of earthquake-induced landslides \citep{lombardo2019geostatistical}, and present a general Bayesian hierarchical modeling framework that has been used in the literature (in various forms and contexts, and with some variations) for jointly modeling multiple occurrences and sizes of various natural phenomena over space \citep{pimont2021prediction,koh2021spatiotemporal,yadav2023joint}. While all (big or small) future landslides matter for appropriate hazard assessment, 
\emph{extreme} landslides are particularly devastating and should therefore be given strong attention in modeling. In the model construction pipeline, we thus advocate using probability distributions motivated by EVT that can adequately capture the upper tail behavior, while simultaneously providing a good fit in the bulk. Specifically, we here review and compare two distinct approaches for modeling the full range of landslide sizes: (i) split (also called `mixture', `spliced', `piecewise', or `hybrid') bulk--tail models \citep{Behrens.etal:2004,Carreau.Bengio:2009,Scarrott.MacDonald:2012,opitz2018inla,castro2019spliced,pimont2021prediction,koh2021spatiotemporal}, wherein the upper tail is modeled with the asymptotically-justified \emph{generalized Pareto (GP)} distribution (see \cite{Davison.Smith:1990,Davison.Huser:2015} and Chapter~2 for more details) and the bulk is modeled using another, essentially arbitrary distribution; and (ii) `subasymptotic' distributions (see, e.g., \cite{Naveau.etal:2016,yadav2021spatial,yadav2021flexible,Stein:2021a,Stein:2021b}, and Chapter~5) capturing the lower tail and the upper tail flexibly in compliance with EVT with a smooth transition in between. While there are several possible subasymptotic distributions, we focus in this chapter on the \emph{extended generalized Pareto (eGP)} distribution \citep{Papastathopoulos.Tawn:2013}, which is a parsimonious parametric model that has found various applications in the study of natural hazards including extreme precipitation \citep{Naveau.etal:2016}, wildfires \citep{cisneros2024deep}, and landslides themselves \citep{yadav2023joint}. To streamline the analysis and the discussion, we here consider only one of the simplest eGP distributions, but see Chapter~5 for more details on other possible options and extensions. From a hazard assessment perspective, both landslide sizes and landslide counts (modeled using the Poisson distribution) are important (see Section~\ref{sec:hazard} for details), and they are here modeled jointly using a latent Gaussian model, in which fixed effects and shared spatially-structured effects can be easily incorporated at the latent level to capture complex spatial trends, spatial and cross-dependencies, as well as covariate effects. Joint modeling of occurrences and sizes is especially key for properly assessing the uncertainty of hazard estimates, which are obtained as a function of both elements.

In this chapter, we showcase the versatility of the proposed general modeling framework using an inventory comprising thousands of landslides simultaneously initiated by the devastating 2008 Wenchuan earthquake in China, going beyond \cite{lombardo2019geostatistical} who studied the spatial distribution of landslides triggered by the same earthquake but did not consider the modeling of landslide sizes. Incorporating physical knowledge into probability models is important to improve both model fit and interpretability, and is one of the recommendations listed in the opinion piece by \cite{huser2024modeling}; using informative geophysical and geomorphological covariates, and defining the model at the `slope unit' resolution (more details in Section~3), are two possible ways to achieve this goal. Therefore, in our case study, both landslides and covariates are observed at the slope unit level, thus requiring a joint areal model for occurrences and sizes, instead of a continuous-space marked point process model. This chapter aims at illustrating this extreme-value-based modeling framework for the study of landslide data, and to compare the pros and cons of split bulk--tail models with subasymptotic models in practice. While we find no uniformly-better approach in this case, we argue that the eGP distribution has an overall better performance and is a more natural modeling solution in general. We also discuss how to perform scalable Bayesian inference in this framework based on a customized efficient Markov chain Monte Carlo (MCMC) algorithm, coupled with latent random effects with a sparse precision matrix (here, chosen with an intrinsic conditional autoregressive (iCAR) probabilistic structure; see \cite{besag1974spatial}). The selection of highly informative covariates (e.g., peak ground acceleration) that can potentially replace the use of latent spatial effects is also discussed, and various latent model structures are compared.
 
The rest of this chapter is structured as follows: Section~\ref{sec:hazard} gives a more precise definition of the `hazard' according to the geoscience literature. Section~3 provides details on the Wenchuan landslide data inventory used in our case study. Section~4 presents the modeling framework and MCMC-based Bayesian inference. Section~5 discusses results for the Wenchuan data application. Finally, Section~6 concludes with a summary of some key points.

\section{Hazard definition}\label{sec:hazard}
Although processes such as floods, earthquakes, or landslides largely differ in their physical manifestation, geophysical and statistical properties, and their impacts, international guidelines on natural hazard prediction share some core requirements \citep{mitchell1993natural}. Broadly speaking, hazard prediction should reflect \emph{where} and \emph{when} (or how frequently) a future natural hazard might occur, and if it does, \emph{how big} (or how destructive) it might be. We also stress that in the geoscience literature, `risk assessment' has a specific meaning that goes beyond `hazard assessment' by additionally encapsulating the \emph{exposure} and \emph{vulnerability} of infrastructure and/or people at risk, and estimating the associated economic, societal and/or environmental costs. In this chapter, we thus distinguish these different notions and adopt the terminology commonly used in geoscience.

The first requirement (`where') involves the notion of \emph{susceptibility}, which indicates how likely a given portion of the landscape is to undergo a given hazard \citep{karlsson2017natural,nicu2022multi}. Statistical models used to address this question in the literature include logistic regression models for presence-absence data, Poisson regression models for count data, or more advanced point process models such as log-Gaussian Cox processes (see, e.g., \cite{lombardo2018point,lombardo2020space}). 

The second requirement (`when' or `how frequently') relates to the \emph{return period} of a given hazard and reflects its recurrence over time. The frequency of such devastating phenomena is often difficult to model because temporal replicates are rarely available and rich spatio-temporal inventories are scarce. Moreover, landslides often occur as part of a sequential compound extreme event (e.g., after a major earthquake or heavy rain), which also complicates the estimation of their return period as this requires an understanding of the frequency of the main triggering factor(s). This aspect of hazard assessment is, therefore, often the most neglected one (but see \cite{dahal2024junction} for a recent attempt to take it into account in a study of rainfall-induced landslides). 

The third requirement (`how big' or `how destructive') involves the notion of \emph{intensity} of an event, which indicates the energy and level of threat associated with a given hazard when it occurs \citep{peng2005space,hungr2018some}. 
Note this notion of intensity differs from that classically used in statistics with point process models. 

In the context of landslides, David J.~Varnes and the International Association of Engineering Geology, Commission on Landslides and Other Mass Movements on Slopes defined the landslide hazard more specifically through statistical terms as \textit{``the probability of occurrence within a specified period and a given area of a potentially damaging phenomenon''} \citep{varnes1984}. This definition was later updated by \cite{guzzetti1999landslidehazard} to explicitly include the intensity of the event. While this general definition was proposed in the context of landslides, the same formulation can be interpreted more broadly and applied similarly in other contexts with various types of natural hazards.

In Section~4, we present an extreme-value-based latent Gaussian modeling pipeline that complies with the first and third requirements of the hazard definition. That is, the proposed model can be used to jointly estimate both the susceptibility and the intensity of a natural phenomenon---here, in the context of earthquake-induced landslides. As there are different ways to characterize the intensity of a natural phenomenon, there is no clear consensus in the literature on how to represent and model it. For instance, flood intensity \citep{vojtek2016flood} is usually expressed as a function of the water height (e.g., \cite{van2023breakthrough}) or peak flow (the maximum rate of discharge; \cite{formetta2021assessment}). 
Earthquake intensity \citep{hough2014earthquake} is commonly expressed as the peak of either the displacement \citep{trugman2019peak}, velocity \citep{dahal2023ground}, or acceleration \citep{murphy1977correlation} experienced at a given location. 
Similarly, landslide intensity \citep{lari2014probabilistic} can also be expressed in multiple ways: for instance, the velocity \citep{he2023modelling} or kinetic energy \citep{pudasaini2021mechanics} associated with a failing mass, the force that the landslide body may exert onto an object \citep{tang2014novel}, its volume \citep{jaboyedoff2020review}, planimetric area \citep{DiNapoli2023} or other shape indices \citep{rana2023landslide}, are all accepted as valid ways to describe the landslide intensity. 
However, most of these elements cannot be measured in the context of large landslide populations. 
Velocities over large landscapes are measurable through Interferometric Synthetic Aperture Radar (InSAR) but only for slow movements \citep{Ahmad2023}. 
Kinetic energy and force can only be estimated through demanding numerical simulations \citep{pudasaini2022landslide}, and their real measurement is conditional on the availability of very expensive instruments installed on a slope for monitoring purposes \citep{mazzanti2015new}. 
Similarly, landslide volumes cannot be measured unless topographic data are available before and after the failure \citep{tseng2013application}. 
Among these landslide intensity parameters, the landslide area (or derivatives thereof) is the easiest one to collect because it is a byproduct of any rigorous landslide mapping procedure \citep{lombardo2021landslide}.    
Hence, out of all the metrics listed above, the landslide area is the most relevant one to support  statistical analyses. In the case study presented in this chapter, we model the square root of the landslide area (easily measurable through remote sensing), which roughly describes the `diameter' of the affected area and can be understood as a proxy for the landslide size and its destructiveness.

\baselineskip 12pt
\bibliographystyle{CUP}
\bibliography{ref}

\end{document}